Jack can paint his apartment in 12 hours. His wife Cheryl requires 20 hours to do the same job. How long will it take them to complete the job if they worked together?
Answer by stanbon(75887)

): Jack can paint his apartment in 12 hours. His wife Cheryl requires 20 hours to do the same job. How long will it take them to complete the job if they worked together?
-----------------------------------
Jack rate = 1/12 job/hr
Cheryl rate = 1/20 job/hr
Together rate = 1/x job/hr
----------------------------------
Equation:
rate + rate = together rate
1/12 + 1/20 = 1/x
Multiply thru by 60x and solve for "x":
---
5x + 3x = 60
8x = 60
x = 7.5 hrs (time required to do the job together)
=====================================================
Cheers,
Stan H.